# Prompt Perturbations Reveal Human-Like Biases in LLM Survey Responses

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used as proxies for human subjects in social science surveys, but their reliability and susceptibility to known response biases are poorly understood. This paper investigates the response robustness of LLMs in normative survey contexts—we test nine diverse LLMs on questions from the World Values Survey (WVS), applying a comprehensive set of 11 perturbations to both question phrasing and answer option structure, resulting in over 167,000 simulated interviews. In doing so, we not only reveal LLMs’ vulnerabilities to perturbations but also show that all tested models exhibit a consistent, but variably intense, recency bias, disproportionately favoring the last-presented answer option. While larger models are generally more robust, all models remain sensitive to semantic variations like paraphrasing and to combined perturbations. By applying a set of perturbations, we reveal that LLMs partially align with survey response biases identified in humans. This underscores the critical importance of prompt design and robustness testing when using LLMs to generate synthetic survey data.
@@
in LLM Survey Responses
Jens Rupprecht1, Georg Ahnert1, and Markus Strohmaier1,2,3 1University of Mannheim, Mannheim 2GESIS - Leibniz Institute for the Social Sciences, Cologne 3Complexity Science Hub, Vienna
@@ 1 Introduction
Large Language Models (LLMs) are increasingly being used as proxies for human subjects in social science research, particularly for generating synthetic responses to survey questions (Argyle et al., 2023; Bisbee et al., 2024, inter alia). This application holds promise for augmenting or replacing costly human data collection, but the reliability of these synthetic respondents and to what extent they overlap with human responses remain open questions. In particular, survey methodology research has found that human responses are sensitive to subtle variations in question and answer phrasing that lead to well-known response biases Krosnick (1991) and it remains unclear whether LLMs, trained on vast amounts of human text, exhibit the same vulnerabilities.
@@
In this paper we present a large-scale empirical study where we investigate the response behavior and robustness of nine different LLMs to normative questions derived from the World Value Survey (WVS; Haerpfer et al., 2022). We systematically apply eleven perturbations, targeting both the question phrasing and the structure of the answer options, as shown in Figure 1. Our research questions are:
@@
We find a consistent recency bias across all tested models in different strengths, where the last-presented answer option is disproportionately favored. Larger models generally exhibit more stable response patterns, but even the largest models are sensitive to changes in question phrasing.
@@
We conduct an experimental evaluation with over 167,000 interviews across nine distinct LLMs, varying in size and origin. We develop and apply a comprehensive set of 11 perturbations, targeting classic survey biases (e.g., response order, scale structure) with common textual variations (e.g., typos, paraphrasing). This makes it a useful baseline to test newly developed LLMs on the same task to put the results into perspective. We provide a detailed analysis of LLM response stability, showing that while some models are more robust than others, all are susceptible to specific types of perturbations. This work underscores the importance of careful prompt and Q&A design when using LLMs as a resource for synthetic survey responses, and provides a framework for evaluating their robustness in survey contexts. [^1] The material can be accessed at https://github.com/dess-mannheim/Survey-Response-Biases-in-LLMs.git.
Figure 1: The Interview Process. The figure shows an example of a bias perturbation (e.g. reversed option order) and a non-bias perturbation (e.g. typos in the question). Each model is interviewed 25 times with each combination.
@@ 2 Related Work
Recent studies explored LLMs as substitutes for human survey participants to generate synthetic data. They found that LLMs can replicate average public opinion on political topics, but often with less variance than human samples (Argyle et al., 2023; Bisbee et al., 2024; Von Der Heyde et al., 2025, inter alia). Others have found that LLM responses can be sensitive to prompting, revealing cultural and demographic biases Geng et al. (2024).
Our work is related to that of Tjuatja et al. (2024), who were among the first to systematically explore human-like response biases in LLMs. They investigated acquiescence, response order, opinion floating, and scale structure effects. Our study extends their work by: (1) using a different, globally diverse survey (the World Values Survey); (2) testing a wider range of LLMs of varying sizes and developers; and (3) incorporating a broader set of perturbations on both answer and question phrasing, such as keyboard typos, paraphrasing, synonyms, priming as well as a combined interaction of two perturbations.
@@ 3 Methods
First, we select a subset of 62 questions representing a sample of different thematic categories, each question in the category sharing the same answer options. These normative, value-oriented Q&A pairs are taken from the WVS’s core variables Haerpfer et al. (2022), excluding all sociodemographic variables. On each Q&A pair we perform all eleven perturbations.
Figure 1 illustrates two exemplary perturbations and the interview process. In total, we perform five perturbations on the answer options as well as five perturbations on the question phrasing of the chosen subset questionnaire. We further include one interaction of two perturbations - one on the answer option and one on the question.
Interviews with the original and each perturbed Q&A pair are carried out 25 times with nine different LLMs. In total, we conducted 167,400 interviews, 18,600 with each model. We compare the distributions of the responded labels for all perturbations through entropy and KL-divergence to the baseline response distribution to the original Q&A pairs. Primacy bias is further examined by comparing the response frequencies of the first and last answer option in the list, whereas opinion floating bias and central tendency bias are tested by checking the shift of responses toward or away from the center of the answer option scale.
@@ 3.1 Experimental Setup
The questions and answer options are sourced from the WVS Wave 7 (2017-2022), a comprehensive cross-national survey on human beliefs and values Haerpfer et al. (2022). From the 259 core variables concerning values and norms, we selected a subset of 62 questions. This subset represents 10 distinct thematic categories, including Trust in People, Confidence in Institutions, Moral Justifiability, and Perception of Democracy, ensuring a diverse range of topics and answer scale formats (e.g., 3-point to 10-point scales).
@@
To ensure our findings are not specific to a single model architecture or developer, we selected nine instruction-tuned LLMs, varying in size, developer, and origin. This selection aims to establish a degree of external validity for our results and includes proprietary and open-source LLMs. The following models were interviewed: Gemini-1.5-Pro, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Llama-3.2-1B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3.5-mini-instruct, Qwen2.5-7B-Instruct and Yi-1.5-6B-Chat.
Infrastructure
Experiments were carried out on a high-performance computing cluster (bwUniCluster 2.0) and a local server equipped with NVIDIA H100 (80GB) GPUs. The total runtime on bwUniCluster 2.0 for one model’s 18,600 interviews, e.g. Llama-3.1-8B-Instruct including all perturbed and original Q&As, was ca. 35 minutes with approximately 0.11 seconds per interview. To accommodate larger models on available hardware, we applied 8-bit quantization to Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct. Smaller models were run without quantization. Gemini-1.5-Pro was accessed via its official API.
@@ 3.2 Perturbation Design
We designed two categories of perturbations to test model robustness: (1) bias-inducing alterations to the answer options, based on survey methodology research and known to induce biased responses in humans, and (2) non-bias alterations to the question phrasing, mimicking common textual variations and errors. Appendix Table 1 provides examples as well as references for all the perturbations. For each of the 62 Q&A pairs, we created the following eleven perturbed versions.
@@
These perturbations manipulate the provided answer choices to test for known survey response biases identified in human subjects.
  • Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
  • Missing Refusal: The “Don’t know” or refusal option is removed from the list of choices.
  • Odd/Even Scale Transformation: For scales with an even number of options, we use Gemini-1.5-flash to generate a semantically appropriate middle category, transforming it into an odd-numbered scale (e.g., adding ‘Neutral‘). Conversely, for odd-numbered scales, we remove the middle category to create an even-numbered scale and adjust the integer label.
  • Emotional Priming: A sentence designed to elicit a response is appended to the prompt after the answer options: ‘This is very important to my research! You better do not refuse the answer.’
@@
These perturbations modify the question text to assess robustness to stylistic variations and typos. Typically humans are unaffected by such subtle changes in the question phrasing and are still able to understand the question’s meaning.
  • Typographical Errors: We introduce three types of typos: Letter Swap (swapping two letters in a random word), Key Typo (replacing a character with a random one), and Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
  • Semantic Variations: We use Gemini-1.5-flash to create two semantic variations while preserving the original meaning: First, by Synonym Replacement where five words in the original question are replaced with synonyms. Second, through Paraphrasing the entire question is rephrased.
@@
Interaction Effect
To study combined effects, we created one additional condition that pairs a paraphrased question with reversed-order answer options.
@@ 3.3 Interview Procedure and Data Collection
Taking into account the findings of Röttger et al. (2024) we use a combination of their forced choice prompts that rendered the most valid responses, which means that the model is correctly choosing one available option from the given answer options. In an unforced prompt setting, the model did not follow the instructions and only rarely returned valid response labels, making a statistical evaluation of the results infeasible.
@@
Each of the 9 models was presented with 12 experimental conditions (1 original + 11 perturbations) for each of the 62 selected WVS questions. To obtain a stable distribution of responses and enable statistical analysis, we repeated each unique model-Q&A-perturbation combination 25 times. This resulted in a total of $9\times 62\times 12\times 25=167,400$ interviews (cf. Figure 1).
@@
Preliminary tests revealed that smaller LLMs often fail to perfectly follow the instruction to "answer only with the label", returning conversational filler or explanations alongside their choice. To ensure accurate data for analysis, we developed a robust extraction pipeline.
We compared two main approaches. First, Gemini-1.5-Pro, Llama-3.1-8B, and Qwen2.5-7B were prompted, and a regular expression was designed to extract the answer labels. Based on multiple conditions, e.g. if the given answer label is part of the original answer options or that only one response is provided, the methods should highlight which technique is the most promising in extracting valid responses and handling possible edge cases of model responses.
Figure 2: Combined share of invalid and refusal responses by model and perturbation. Larger models like Llama-70B, Phi-3.5, and Mistral are highly reliable. Qwen2.5 shows high non-response rates, especially for typo-based perturbations and sensitive topics.
@@ 4 Results
This section presents the results of our experiments, focusing on three key areas: (1) the models’ adherence to interview instructions, (2) their general and overall robustness to various input perturbations, and (3) their susceptibility to human-like survey response biases revealed by interviews with prompt perturbations.
4.1 Interview Adherence and Refusal Rates
Notably, we observed topic-specific sensitivity. For questions regarding the Perception of Elections, Qwen2.5-7B failed to provide a valid, on-scale answer in 91.3% of cases, even for the original, unperturbed questions. This might suggest the presence of strong content-based guardrails or restrictions in certain models (cf. Figure 8).
@@ 4.1 Interview Adherence and Refusal Rates
4.2 Consistency and Robustness to Question and Answer Perturbations
We distinguish between response consistency (the tendency of a model to give the same answer to the same prompt, measured by entropy) and response robustness (the tendency to maintain a similar answer distribution under perturbation, measured by KL-divergence). A KL divergence of zero indicates a perfect match and thus full robustness against the input perturbation, whereas a high entropy value indicates very inconsistent response behavior.
Effect of Model Size and Consistency
First, we found that model size is in an inverse relationship with response consistency; smaller models exhibited higher entropy and standard deviation when asked the same question multiple times, indicating more random response behavior.
When assessing robustness to perturbations, we found a clear relationship with model size: larger models tend to be more robust. Figure 4 shows the percentage of questions for which the models produced a perfectly identical response distribution (KL-Div = 0) despite perturbations. Llama-3.3-70B and Gemini-1.5-Pro were the most robust, often replicating their original answers in over 50% of cases. The smaller Llama models were the least robust, with Llama-3.2-1B perfectly replicating its answers in fewer than 5% of cases on average. This suggests that scale is a key factor in achieving stable response behavior in synthetic response generation.
Figure 4: Share of fully robust responses (KL-Divergence = 0) by model and perturbation type. Top: Non-bias perturbations. Bottom: Bias perturbations. A clear trend emerges: larger models (Llama-70B, Gemini) are substantially more robust than smaller models (Llama-1B, Llama-3B). Certain perturbations, like the combined interaction effect, paraphrasing, and reversing the scale, are challenging for all models.
Effect of Perturbation Type
Further, Figure 4 highlights the share of fully robust responses (Kullback-Leibler Divergence = 0) across all questions by perturbation and LLM. It shows that certain perturbations had a greater impact on robustness across all models.

### v3

# Prompt Perturbations Reveal Human-Like Biases in Large Language Model Survey Responses

@@ Abstract
Large Language Models (LLMs) are increasingly used as proxies for human subjects in social science surveys, but their reliability and susceptibility to known human-like response biases, such as central tendency, opinion floating and primacy bias are poorly understood. This work investigates the response robustness of LLMs in normative survey contexts—we test nine LLMs on questions from the World Values Survey (WVS), applying a comprehensive set of ten perturbations to both question phrasing and answer option structure, resulting in over 167,000 simulated survey interviews. In doing so, we not only reveal LLMs’ vulnerabilities to perturbations but also show that all tested models exhibit a consistent recency bias, disproportionately favoring the last-presented answer option. While larger models are generally more robust, all models remain sensitive to semantic variations like paraphrasing and to combined perturbations. This underscores the critical importance of prompt design and robustness testing when using LLMs to generate synthetic survey data.
@@
in Large Language Model Survey Responses
Jens Rupprecht1, Georg Ahnert1, and Markus Strohmaier1,2,3 1University of Mannheim, Mannheim 2GESIS – Leibniz Institute for the Social Sciences, Cologne 3Complexity Science Hub, Vienna
@@ 1 Introduction
Large Language Models (LLMs) are increasingly being used as proxies for human subjects in social science research, particularly to generate synthetic responses to survey questions (Argyle et al., 2023; Bisbee et al., 2024, inter alia). This application holds promise in augmenting or replacing costly human data collection, but the reliability of these synthetic respondents and to the extent of overlap with human responses and response biases remain open questions. In particular, research in survey methodology has found that human responses are sensitive to subtle variations in question and answer phrasing that lead to well-known response biases Krosnick (1991) and it remains unclear whether LLMs, trained on vast amounts of human text, exhibit the same vulnerabilities.
@@
We present a large-scale empirical study, with 167,000 survey interviews in total, investigating the response behavior and robustness of nine different LLMs to normative questions derived from the World Value Survey (WVS; Haerpfer et al., 2022). By developing and applying a comprehensive set of ten perturbations both in the structure of the answer options (Table 1), such as typos or synonyms, and in the question phrasing (Table 2), such as e.g., changes to response order or scale structure, we answer the following research questions.
@@
Next to the perturbation framework, we provide a detailed analysis of LLM response robustness, showing that while some models are more robust than others (e.g. Llama-3.3-70B-Instruct and Gemini-1.5-Pro), all are susceptible to specific types of perturbations. We find a consistent recency bias across all tested models in different strengths, where the last-presented answer option is disproportionately favored up to 20 times. Larger models generally exhibit more robust response patterns, but even the largest models are sensitive to changes in question phrasing. Thus, this work underscores the importance of careful prompt and Q&A design when using LLMs as a resource for synthetic survey responses. The perturbation framework can serve as a useful baseline to test newly developed LLMs to put the results into perspective and to generally evaluate LLM robustness in survey contexts. Further, we intend to make the Q&A dataset with perturbations available for other researchers to benchmark and check response biases in newer or other LLMs.
Figure 1: The Interview Process. The figure displays an example of a answer option perturbation (a bias perturbation, e.g. reversed option order) and an question perturbation (a non-bias perturbation, e.g. typos in the question). Each model is prompted 25 times with every perturbation as well as the original Q&A phrasing. All responses are collected, processed and statistically analyzed.
@@ 2 Related Work
Recent studies explored LLMs as substitutes for human survey participants to generate synthetic data. They found that LLMs can replicate average public opinion on political topics, but often with less variance than human samples (Argyle et al., 2023; Bisbee et al., 2024; von der Heyde et al., 2025; Dominguez-Olmedo et al., 2024, inter alia). Others have found that LLM responses can be sensitive to prompting, revealing cultural and demographic biases Geng et al. (2024).
Our work is related to that of Tjuatja et al. (2024), who were among the first to systematically explore human-like response biases in LLMs. They investigated acquiescence, response order, opinion floating, and scale structure effects. Our study extends their work by: (1) using a different, globally diverse survey (the World Values Survey); (2) testing a wider range of LLMs; and (3) incorporating a broader set of perturbations on both answer and question phrasing, such as keyboard typos, paraphrasing, synonyms, priming as well as a combined interaction of two perturbations.
@@ 3 Methods
First, we select a subset of 62 questions representing a sample of different thematic categories, each question in the category sharing the same answer options. These normative, value-oriented Q&A pairs are taken from the WVS’s core variables Haerpfer et al. (2022), excluding all sociodemographic variables. Second, we perform the ten perturbations mentioned in Section 3.2 each Q&A pair.
Figure 1 illustrates two exemplary perturbations and the interview process. In total, we perform five perturbations on the answer options as well as five perturbations on the question phrasing of the chosen subset questionnaire. We further include one interaction of two perturbations, one on the answer option and one on the question.
Third, we carry out interviews with the original and each perturbed Q&A pair 25 times with nine different LLMs. In total, we conducted 167,400 interviews, 18,600 with each model. Last, we compare the distributions of the responded labels for all perturbations through entropy and Kullback-Leibler divergence (KL divergence) to the baseline response distribution to the original Q&A pairs. Primacy bias is further examined by comparing the response frequencies of the first and last answer option in the list, whereas opinion floating bias and central tendency bias are tested by checking the shift of responses toward or away from the center of the answer option scale.
@@ 3.1 Experimental Setup
The questions and answer options are sourced from the WVS Wave 7 (2017-2022), a comprehensive cross-national survey on human beliefs and values Haerpfer et al. (2022). The 259 core WVS Q&A pairs represent 10 distinct thematic categories, including Trust in People, Confidence in Institutions, Moral Justifiability, and Perception of Democracy, ensuring a diverse range of topics and answer scale formats (e.g., 3-point to 10-point scales). We used stratified sampling to select six to seven Q&A pairs per thematic category, resulting in a total of 62 Q&A pairs.
@@
To ensure our findings are not specific to a single model architecture or developer, we selected nine instruction-tuned LLMs, varying in size, developer, and origin. This selection aims to establish a degree of external validity for our results and includes proprietary and open-source LLMs. The following models were interviewed: Gemini-1.5-Pro Georgiev et al. (2024), Llama-3.3-70B-Instruct Meta (2025b), Llama-3.1-8B-Instruct Grattafiori et al. (2024), Llama-3.2-3B-Instruct, Llama-3.2-1B-Instruct Meta (2025a), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), Phi-3.5-mini-instruct Abdin et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2025) and Yi-1.5-6B-Chat Young et al. (2024).
@@ 3.2 Perturbation Design
We designed two categories of perturbations to test model robustness: (1) bias-inducing alterations to the answer options, based on survey methodology research and known to induce biased responses in humans Tjuatja et al. (2024), and (2) non-bias alterations to the question phrasing, mimicking common textual variations and errors. Table 1 provides examples as well as references for all the perturbations. For each of the 62 Q&A pairs, we created the following ten perturbed versions.
Type Perturbation Question Answer Options Bias and Reference
Original
Original
For each of the following aspects,
indicate how important
it is in your life.
Would you say it is very important,
rather important, not very important
or not important at all?
Family
[’1=Very important ’, ’2=Rather important ’,
’3=Not very important ’, ’4=Not important at all’,
"-1=Don’t know"]
Haerpfer et al. (2022)
@@
(1) Reversed Response Order
For each of the following aspects,
indicate how important
it is in your life.
Would you say it is
very important,
rather important,
not very important
or not important at all?
Family
["-1=Don’t know", ’4=Not important at all’,
’3=Not very important ’, ’2=Rather important ’,
’1=Very important ’]
Primacy Bias
Tjuatja et al. (2024); Krosnick and Alwin (1987); Kampen (2007); O’Halloran et al. (2014)
(2) Missing Refusal Option
[’1=Very important ’, ’2=Rather important ’,
’3=Not very important ’, ’4=Not important at all’]
Opinion Floating Bias
Schuman and Presser (2000); Tjuatja et al. (2024)
(3) Odd/Even Scale Transformation
[’1=Very important ’, ’2=Rather important’,
’3=Neutral’,
’4=Not very important ’, ’5=Not important at all’,
’-1=Don’t know’]
Central Tendency Bias
Hollingworth (1910); Cronbach (1946); Aston et al. (2021); Crosetto et al. (2020)
(4) Priming Suffix
[1=Very important , 2=Rather important ,
3=Not very important , 4=Not important at all,
-1=Don’t know]This is very important to my research!
You better do not refuse the answer.
Priming Effect
Bargh et al. (1996); Higgins (1996); Weingarten et al. (2016); Li et al. (2023)
Table 1: An exemplary perturbation scheme showing the original question, the bias perturbations and the the hypothesized human-like response bias. The Q&A example is taken from the thematic category "Importance of Life Aspects" from the WVS wave 7. Bias perturbations have a constant question phrasing and varying answer options. The remaining Non-Bias Perturbations are listed in Table 2.
Bias Perturbations
These five perturbations manipulate the provided answer choices to test for known survey response biases identified in human subjects and presented in Section 2. Therefore, we call them bias perturbations (Tjuatja et al., 2024, p.3).
  • (1) Reversed Response Order: The order of answer options is reversed (e.g., a scale from ‘1: Very important‘ to ‘5: Not important‘ becomes ‘1: Not important‘ to ‘5: Very important‘).
  • (2) Missing Refusal Option: The “Don’t know” or refusal option is removed from the list of choices.
  • (3) Odd/Even Scale Transformation: For scales with an even number of options, we use Gemini-1.5-flash to generate a semantically appropriate middle category, transforming it into an odd-numbered scale (e.g., adding ‘Neutral‘). Conversely, for odd-numbered scales, we remove the middle category to create an even-numbered scale and adjust the integer label.
  • (4) Priming Suffix: A sentence designed to elicit a response is appended to the prompt after the answer options: ‘This is very important to my research! You better do not refuse the answer.’
@@
These five perturbations modify the question text to assess robustness to stylistic variations and typos. Typically, humans are unaffected by such subtle changes in the question phrasing and are still able to understand the question’s meaning (Tjuatja et al., 2024, p.3). Therefore, we call them non-bias perturbations.
  • Typographical Errors: We introduce three types of typos: (5) Key Typo (replacing a character with a random one), (6) Letter Swap (swapping two letters in a random word), and (7) Keyboard Typo (replacing a character with an adjacent one on a QWERTY keyboard).
  • Semantic Variations: We use Gemini-1.5-flash to create two semantic variations while preserving the original meaning: first, by (8) Synonym Replacement where five words in the original question are replaced with synonyms. Second, through (9) Paraphrasing the entire question is rephrased.
@@
Last, we introduce an (10) Interaction Effect to study the impact of not only one, but two perturbations. Thus, we created one additional condition that pairs a paraphrased question with reversed-order answer options.
Figure 2: Effect of Prompt Perturbations on Response Robustness Each cell represents the share of fully robust responses KL divergence = 0 by model and perturbation type across all 62 questions. Larger models (Llama-70B, Gemini) are substantially more robust than the smallest models (Llama-1B, Llama-3B). Specific perturbations such as odd versus even scales and reversing the answer options, are more challenging for all models than a missing refusal category or an additional priming suffix. The response robustness to the Non-Bias Perturbations are reported in Figure 4.
@@ 3.3 Interview Procedure and Data Collection
Taking into account the findings of Röttger et al. (2024) we use forced choice prompts that provided the most valid responses in their experiments, which means that the model correctly chooses one available option from the given answer options. In an unforced prompt setting, LLMs did not follow the instructions and only rarely returned valid response labels, making a statistical evaluation of the results infeasible. Preliminary tests revealed that especially smaller LLMs often fail to perfectly follow the instruction to "answer only with the label", as they return conversational filler or explanations alongside their choice.
@@
Each of the 9 models was presented with 12 experimental conditions (1 original + ten perturbations, where for the perturbation Odd/Even Scale Transformation one run was done for the odd and even scale scenario) for each of the 62 selected WVS questions. To obtain a stable distribution of responses and enable statistical analysis, we repeated each unique model-Q&A-perturbation combination 25 times. This resulted in a total of $9\times 62\times 12\times 25=167,400$ interviews.
@@
To ensure accurate data for analysis, we developed a robust extraction pipeline. We compared two main approaches. First, Gemini-1.5-Pro, Llama-3.1-8B, and Qwen2.5-7B were prompted, and a regular expression was designed to extract the answer labels. Based on multiple conditions, e.g. if the given answer label is part of the original answer options or that only one response is provided, the methods should highlight which technique is the most promising in extracting valid responses and handling possible edge cases of model responses.
@@ 4 Results
This section presents the results of our experiments, focusing on two key research questions: (1) LLMs general and overall robustness to various input perturbations, (2) their susceptibility to human-like survey response biases revealed by interviews with bias prompt perturbations. Further, we also investigate the models’ general adherence to interview instructions.
@@
4.1 Robustness to Question and Answer Perturbations (RQ1)
We distinguish between response robustness (the tendency to maintain a similar answer distribution under perturbation, measured by KL divergence) and response consistency (the tendency of a model to give the same answer to the same prompt, measured by entropy). A KL divergence of zero indicates a perfect match and thus full robustness against the input perturbation, whereas a high entropy value indicates very inconsistent response behavior.
Effect of Model Size on Robustness
First, when assessing robustness to perturbations, we found a clear relationship with model size: larger models tend to be more robust. Figures 2 and 4 show the percentage of questions for which the models produced a perfectly identical response distribution (KL divergence = 0) despite perturbations. Llama-3.3-70B and Gemini-1.5-Pro were the most robust, often replicating their original answers in over 50% of cases. The smaller Llama models were the least robust, with Llama-3.2-1B perfectly replicating its answers in fewer than 5% of cases on average. This suggests that scale is a key factor in achieving stable response behavior in synthetic response generation.
Second, we found that model size is in an inverse relationship with response consistency; smaller models exhibited higher entropy and standard deviation when asked the same question multiple times, indicating more random response behavior (Figure 9).
Effect of Perturbation Type on Robustness
Further, Figures 2 and 4 highlight the share of fully robust responses (KL divergence = 0) across all questions by perturbation and LLM. It shows that some perturbations had a greater impact on robustness across all models.